Marcia Juarez and her brother-in-law have a successful business with several employees. They decide
to borrow $12,000 to pay their next quarterly deposits for payroll tax and federal income tax. They get
the money at 6% compounded monthly and repay all interest and principal after 3 months. How much
do they repay?
1
Expert's answer
2020-12-15T14:11:06-0500

"k=6\%,~n=3,~S_0=\$12000."

"S=S_0(1+\frac{k}{12\cdot100\%})^n=\\=\$12000\cdot1.005^3=\$12180.90" .